# Ekpyrosis in Quantum Gravitational Anisotropic Bouncing Models

### v1

@@ Abstract
We explore the isotropization of the universe starting from potentially large anisotropies in the bouncing models using the ekpyrotic mechanism. As an example of a concrete non-singular bouncing mechanism, we consider the effective description of loop quantum cosmology for Bianchi-I and Bianchi-IX spacetimes for ekpyrotic and ekpyrotic-like potentials. For both of these spacetimes the cosmological singularity is resolved via multiple short-duration non-singular bounces. We perform a large number of numerical simulations for a wide range of initial conditions and find that the relative strength of the anisotropies at the end of the bounce regime is noticeably reduced in more than $90\%$ of the simulations, providing strong evidence for the isotropization ability of the ekpyrotic potentials. While the ekpyrosis phase in all the simulations is found to be rather short-lived, isotropization occurs over cycles of rapid non-singular bounces in the Planck regime via enhancement of the contribution of the (isotropic) energy density relative to the anisotropies at the bounces. Achieving isotropization is found to be easier in Bianchi-I spacetimes when compared to Bianchi-IX spacetimes. Our results demonstrate that, while ekpyrosis might itself be insufficient to tame anisotropies at a single bounce, it can be significant when coupled with non-singular cycles in the bounce regime.
@@ I Introduction
In recent years, various investigations have been carried out on bouncing cosmologies to determine whether they can serve as potentially viable alternatives to inflation. The typical scenario in these models consists of a contracting macroscopic universe that undergoes a bounce, from which an expanding universe emerges. The first challenge in constructing such bouncing models is the cosmic singularity which is inevitable in general relativity (GR) unless the null energy condition is violated which often results in pathologies. For a viable bouncing scenario, it is required that the singularity resolution be a generic phenomena and not a result of fine-tuning of initial conditions or the choice of some exotic matter. If the fundamental challenge of obtaining singularity resolution is addressed, bouncing models face the next challenge – the growth of anisotropies in the contracting branch. As the mean scale factor decreases, the anisotropic shear grows at a faster rate than the energy density of matter source with an equation of state $w<1$. As a result, the approach to big bang/big crunch singularities is dictated by anisotropies for such a matter source resulting in a mixmaster dynamics and BKL instability [1, 2, 3] which can potentially also affect the viability of a non-singular bounce. Even in the simplest anisotropic models such as Bianchi-I, anisotropies dominate the approach to singularity [4]. Thus, even if a bounce occurs, it is expected to be highly anisotropic and the resulting post-bounce universe may also retain this character. The pertinent question is whether one can construct robust non-singular bouncing models in anisotropic spacetimes where anisotropies are tamed in the contracting branch, allowing the universe to isotropize.
@@
To answer this question, one needs two main ingredients: a mechanism to resolve the singularities in the anisotropic setting and a mechanism to diminish the anisotropies. Let us first consider the resolution of singularities. It has long been expected that to robustly address the problem of cosmological singularities, one would need some quantum gravitational input. One of the arenas where this has been rigorously demonstrated is loop quantum cosmology (LQC) [5] where cosmological singularities have been demonstrated to be generically resolved without imposing any fine-tuned initial conditions or any violation of energy conditions [6, 7]. In LQC, the big bang singularity no longer exists and an expanding universe emerges from a contracting universe which experiences a bounce caused by non-perturbative quantum geometry effects which dominate the dynamics when the energy density reaches Planckian values during contraction [8, 9]. This result has been found to be robust for various spacetimes [10], including in the presence of spatial curvature [11, 12, 13], cosmological constant [14, 15], anisotropies [16, 17] and Fock quantized inhomogeneities [18]. The results of singularity resolution have also been extended to black hole spacetimes (see [19] for a review). These results indicate that the problem of singularities is robustly addressed in LQC, and for this reason we will consider this as a concrete example for a non-singular bouncing mechanism in this manuscript.
@@
The advantage of choosing LQC is that a fair amount of work has been done to understand the phenomenology of the anisotropic models in LQC. While the underlying geometry in LQC is discrete and is described by difference equations, for phenomenological explorations it is common to use the continuum effective description of LQC, which is based on differential equations and is shown to faithfully incorporate the underlying quantum geometry corrections in various isotropic and anisotropic spacetimes including the Bianchi-I model [20, 21, 22, 17]. Using effective spacetime description, where one can express quantum gravitational modifications in terms of modifications to the classical Hamiltonian [23, 9], it has been shown that irrespective of the choice of the matter content all strong curvature singularities are generically resolved, and isotropic and anisotropic cosmological spacetimes are geodesically complete [24, 25, 26, 27, 28, 29, 30]. The energy density and anisotropic shear are bounded in Bianchi-I, II and IX spacetimes [31, 32, 33, 34, 35]. The Bianchi-I and Bianchi-IX models in LQC have an especially rich phenomenology of anisotropic Kasner transitions at the bounce [36, 37, 38, 39], allowing us to test our ideas under a variety of initial conditions. It has been demonstrated that a viable non-singular inflationary model can be constructed starting from highly anisotropic initial conditions in Bianchi-I effective spacetime [40, 41].
@@
As far as the issue of suppression of anisotropies is concerned, perhaps the most straightforward way is to consider a source of matter-energy which has an equation of state $w>1$. In this the energy density of matter grows faster than anisotropic shear and it is possible that the universe can isotropize as the scale factor decreases. Such an ultra-stiff equation of state can occur in ekpyrotic scenarios [42, 43]. These models were originally motivated by bulk-brane dynamics which have a negative exponential potential, and as the branes move towards a collision the moduli field has a pressure greater than the energy density and the ekpyrotic field behaves as an isotropic fluid having an equation of state larger than unity. In the ekpyrotic phase, the energy density grows faster than anisotropies as the universe contracts towards the singularity potentially leading to its isotropization and potentially avoiding the BKL instability [44, 45]. Recently, ekpyrotic mechanisms have been used in models with ultra-slow contraction to explain homogeneity, flatness and isotropy in the contracting branch [46, 47]. Models which are ekpyrotic-like, with a different negative potential, can also potentially allow an isotropization [48, 49]. If the ekpyrotic field is combined with a bounce, such as in LQC, one can expect the bounce to isotropize, at least when one starts with anisotropies which are small. While the intuitive reason for the ekpyrotic field to diminish the growth of anisotropies in the contracting branch is easy to understand, whether or not sufficient and effective ekpyrosis occurs depends on the details of the background dynamics and is potentially sensitive to the initial conditions, including the magnitude of initial anisotropic shear compared to matter energy density. Therefore an important question with any ekpyrotic mechanism is its duration and strength and whether it can result in isotropization for arbitrary initial conditions, especially if the anisotropies are large. An example of this occurs in bouncing models in LQC where ekpyrotic potentials have been considered in both isotropic and anisotropic models showing that a non-singular evolution is possible, but a recent study for isotropic model showed that the phase of ekpyrosis though generic, is rather small in duration and weak in amplitude [50]. Answering the above question on the duration and strength of ekpyrosis is also important to construct viable models of the early universe where curvature perturbations start in the contracting branch before the bounce, such as in the matter-bounce alternative to inflation [51, 52, 53, 50].
@@
The goal of this manuscript is to understand whether ekpyrosis can result in an effective isotropization of the non-singular bounce. We consider Bianchi-I and Bianchi-IX models in the effective spacetime description of LQC as our setting and contrast the dynamics when the matter-energy is sourced by a massless scalar field only versus the ekpyrotic field both in the ekpyrotic potential [43] as well as an ekpyrotic-like potential [48]. Specifically, we study whether the ekpyrotic field can dominate the bounce regime and lead to a reduction in the strength of the anisotropic shear relative to the energy density of the universe (which is assumed to be isotropic). Earlier studies with ekpyrotic potential in LQC confirm that non-singular cyclic evolution is possible [54, 55], but it is only in the presence of anisotropies [56] that the moduli field can turn around which is necessary for a viable ekpyrotic scenario as proposed originally [42]. Similarly, attempts have been made to explore the effect of anisotropic evolution on the cosmological perturbations in LQC, however these studies also restrict anisotropies assuming a point-like bounce and a fixed ultra-stiff equation of state [57, 58]. Recall that in general, the anisotropic approach to singularity (and hence the bounce) would be cigar-like with two scale factors contracting and one expanding before the bounce. To obtain a point-like bounce (all scale factors contracting pre-bounce) under generic anisotropic conditions it is necessary to have a viable ekpyrotic scenario. Moreover, in general, the equation of state of the ekpyrotic field varies during the evolution, and previous studies indicate that it becomes ultra-stiff only for a short duration around the bounce. Thus, it is not clear at the outset whether these short phases of ekpyrosis can cause sufficient isotropization. In this manuscript, we go beyond the assumption of small anisotropic perturbations and explore whether an isotropic universe can be obtained starting from generic initial conditions, considering the universe to be fully anisotropic initially, in the presence of an ekpyrotic potential without making restrictions on the equation of state.
@@
The effective equations of motion for the loop quantized Bianchi-I and Bianchi-IX spacetimes are very complicated and efforts are still underway to develop an intuitive understanding of the interplay between the energy density and the anisotropic shear in these models. Recently, a parabolic relation between the anisotropic shear and the energy density has been found to hold at the bounce for certain matter fields in the Bianchi-I model using extensive numerical simulations [59]. This is very surprising in light of the complicated nature of the effective equations and the fact that no such simple relation is found to hold before or after the bounce. However, the current state of the art is far from the level of analytical control needed to go beyond making a few qualitative remarks, and the progress is more often based on numerical simulations. Consequently, our analysis in this manuscript is numerical, yet wide enough to gain general insights on possible outcomes from a matter-ekpyrotic bounce. We consider the effective Bianchi-I and Bianchi-IX spacetimes for ekpyrotic and ekpyrotic-like potentials. We consider two different values of the strengths of the potential and perform $500$ numerical simulations with varying initial conditions chosen randomly in regime when the universe is macroscopic and contracting towards a bounce. To measure the extent of the isotropization achieved due to ekpyrotic potentials, we use the massless scalar field as the reference case against which the extent of isotropization is measured. Specifically, each simulation with the ekpyrotic or ekpyrotic-like potentials is accompanied by a corresponding simulation with only the massless scalar field with same initial conditions for three triads, two connection components and energy density. Comparing the relative strength of the anisotropies at the bounce (relative to the matter energy density) we state that isotropization has been achieved if the relative strength of the anisotropies has decreased in comparison with the simulation with massless scalar field with matching initial conditions.
@@
From about 4000 simulations performed in this work, we find that while there are small differences in the effectiveness of the two different ekpyrotic potentials considered, isotropization is achieved in an overwhelming majority of the simulations, in both the Bianchi-I and Bianchi-IX models when the ekpyrotic field is present. Further, the equation of state of the ekpyrotic fields becomes ultra-stiff only for a small duration near the bounce, which by itself may not be sufficient to produce noticeable isotropization. However, we found that with the ekpyrotic fields as the matter content, a contracting universe generically undergoes multiple non-singular bounces in rapid succession before the emergence of a macroscopic expanding universe, causing multiple short spurts of ekpyrosis. These short spurts of ekpyrosis coupled with multiple bounces is found to be sufficient to produce isotropization. Another important insight we obtain is that the effect of the ekpyrotic potential on the anisotropic shear is actually indirect. This information is distilled by analyzing the effect of the ekpyrotic field on the energy density and the anisotropic shear, separately, in each simulation. We find that the ekpyrotic potential leads to a direct rise in the energy density at the bounce, indicated by a rise in the energy density at the bounce compared to the massless scalar field case in an overwhelming majority of the simulations. However, the effect of ekpyrosis on the anisotropic shear is not as consistent. Since the contribution of the energy density at the bounce is increased in almost all cases, the relative strength of the anisotropies goes down in almost all cases. It turns out that the isotropization is achieved in a slightly higher fraction of cases in the Bianchi-I model as compared to the Bianchi-IX model. Further, we find that increasing the strength of the ekpyrotic potentials does result in a higher fraction of cases where isotropization is achieved. These findings are summarized in a table at the end of the manuscript which provides a more granular look at these results.

### v2

@@ Abstract
We explore the isotropization of a model anisotropic universe in the bouncing models using the ekpyrotic potential without assuming initial conditions corresponding to an ekpyrotic phase. In particular, we explore the way use of ekpyrotic potentials may dynamically help isotropization for the considered initial conditions corresponding to the macroscopic classical contracting universe with potentially large anisotropies. As an example of a concrete non-singular bouncing mechanism, we consider the effective description of loop quantum cosmology for Bianchi-I and Bianchi-IX spacetimes for ekpyrotic and ekpyrotic-like potentials. Considering two different values of potential parameters and initial conditions corresponding to a classical macroscopic universe, we show that for both of these spacetimes the cosmological singularity is resolved via multiple short-duration non-singular bounces caused by quantum gravitational effects. We perform a large number of numerical simulations for a wide range of initial conditions which do not favor ekpyrosis initially. Even with such unfavorable initial conditions we show that the relative strength of the anisotropies at the end of the bounce regime is noticeably reduced in more than $90\%$ of the simulations.This provides a strong evidence for the isotropization ability of the ekpyrotic potentials. We find that isotropization can occur over cycles of rapid non-singular bounces in the Planck regime via enhancement of the contribution of the (isotropic) energy density relative to the anisotropies at the bounces. Achieving isotropization is found to be easier in Bianchi-I spacetimes when compared to Bianchi-IX spacetimes. Our results demonstrate that, even with initial conditions which are not most favorable for existence of ekpyrosis, an effective isotropization can occur in non-singular anisotropic models with ekpyrotic and ekpyrotic-like potentials.
@@ I Introduction
In recent years, various investigations have been carried out on bouncing cosmologies to determine whether they can serve as potentially viable alternatives to inflation. A typical scenario in these models consists of a contracting macroscopic universe from the classical regime that undergoes a bounce, from which an expanding universe emerges. The first challenge in constructing such bouncing models is the cosmic singularity which is inevitable in general relativity (GR) unless the null energy condition is violated which often results in pathologies. For a viable bouncing scenario, it is required that the singularity resolution be a generic phenomena and not a result of fine-tuning of initial conditions or the choice of some exotic matter. If the fundamental challenge of obtaining singularity resolution is addressed, bouncing models face the next challenge – the growth of anisotropies in the contracting branch. As the mean scale factor decreases, the anisotropic shear grows at a faster rate than the energy density of matter source with an equation of state $w<1$. As a result, the approach to big bang/big crunch singularities is dictated by anisotropies for such a matter source resulting in a mixmaster dynamics and BKL instability [1, 2, 3] which can potentially also affect the viability of a non-singular bounce. Even in the simplest anisotropic models such as Bianchi-I spacetime, anisotropies dominate the approach to singularity [4]. Thus, even if a bounce occurs, it is expected to be highly anisotropic and the resulting post-bounce universe may also retain this character. The pertinent question is whether one can construct robust non-singular bouncing models in anisotropic spacetimes where anisotropies are tamed in the contracting branch, allowing the universe to isotropize.
@@
To answer this question, one needs two main ingredients: a mechanism to resolve the singularities in the anisotropic setting and a mechanism to diminish the anisotropies. Let us first consider the resolution of singularities. It has long been expected that to robustly address the problem of cosmological singularities, one would need some quantum gravitational input. One of the arenas where this has been rigorously demonstrated is loop quantum cosmology (LQC) [5] where cosmological singularities have been demonstrated to be generically resolved without imposing any fine-tuned initial conditions or any violation of energy conditions [6, 7]. In LQC, the big bang singularity no longer exists and an expanding universe emerges from a contracting universe which experiences a bounce caused by non-perturbative quantum geometry effects which dominate the dynamics when the energy density reaches Planckian values during contraction [8, 9]. To understand singularity resolution and the underlying physics at the Planck scale in LQC, one considers initial states in a macroscopic classical universe and evolves them towards classical big bang using quantum Hamiltonian constraint. As one approaches, Planck regime there are departures from the classical Einstein’s equations due to quantum gravitational effects which result in a bounce when the spacetime curvature becomes Planckian. This result has been found to be robust for various spacetimes [10], including in the presence of spatial curvature [11, 12, 13], cosmological constant [14, 15], anisotropies [16, 17] and Fock quantized inhomogeneities [18]. The singularity resolution in LQC occurs because of the underlying quantum geometry effects which lead to maximum allowed values of energy density and anisotropic shear [19]. The results of singularity resolution have also been extended to black hole spacetimes (see [20] for a review). These results indicate that the problem of singularities is robustly addressed in LQC, and for this reason we will consider this as a concrete example for a non-singular bouncing mechanism in this manuscript.
@@
The advantage of choosing LQC is that a fair amount of work has been done to understand the phenomenology of the anisotropic models in LQC. While the underlying geometry in LQC is discrete and is described by difference equations, for phenomenological explorations it is common to use the continuum effective description of LQC, which is based on differential equations and is shown to faithfully incorporate the underlying quantum geometry corrections in various isotropic and anisotropic spacetimes including the Bianchi-I model [21, 22, 23, 17]. The validity of effective description is found using extensive numerical simulations where as described above one starts from a macroscopic classical universe. Using effective spacetime description, where one can express quantum gravitational modifications in terms of modifications to the classical Hamiltonian [24, 9], it has been shown that irrespective of the choice of the matter content all strong curvature singularities are generically resolved, and isotropic and anisotropic cosmological spacetimes are geodesically complete [25, 26, 27, 28, 29, 30, 31]. The energy density and anisotropic shear are bounded in Bianchi-I, II and IX spacetimes [32, 19, 33, 34, 35]. The Bianchi-I and Bianchi-IX models in LQC have an especially rich phenomenology of anisotropic Kasner transitions at the bounce [36, 37, 38, 39], whose physical implications have been studied (see for eg. [40, 41, 42, 43]). Note that in our manuscript, we focus on anisotropies in the context of bouncing models in LQC but there have been various other studies in the context of pre-big bang scenarios, see for eg. [44, 45].
As far as the issue of suppression of anisotropies is concerned, perhaps the most straightforward way is to consider a source of matter-energy which an equation of state $w>1$. In this case the energy density of matter can grow faster than anisotropic shear and it is possible that the universe can isotropize as the scale factor decreases. Such an equation of state can occur in ekpyrotic scenarios [46, 47] where it is possible to have $w\gg 1$, a condition necessary for ekpyrosis. In this manuscript, we would label such an equation of state as ultra-stiff. These models were originally motivated by bulk-brane dynamics where inter-brane dynamics is captured by the evolution of moduli field in a negative exponential potential. The dynamics of the ekpyrotic field results in different phases of evolution of the universe. In the ekpyrotic potential (see eq. (3.1)) when the field is in the large positive regime and moving towards smaller $\phi$ values the universe is classical and macroscopic and in a contracting phase. The field then enters a region of negative well in the potential where conditions favorable for ekpyrosis can arise. As the field approaches large negative values, branes come close to each other and eventually collide causing a big crunch. After the collision of branes, the field moves from large negative values towards positive values and the universe enters an expanding phase from a small size to macroscopic values. Subsequently, the field reaches positive values, and the universe becomes macroscopic, Hubble friction stops the growth of the field causing it to turn around and the cycle repeats. Note that as the branes move towards a collision (or come out of a collision), the moduli field can have a pressure greater than the energy density and the ekpyrotic field behaves as an isotropic fluid having an equation of state larger than unity. Our goal in this manuscript is to study the evolution of macroscopic contracting universe starting with large anisotropies in ekpyrotic and ekpyrotic-like potentials without assuming any conditions which may initially favor or guarantee ekpyrosis. In particular, we consider initial conditions corresponding to the classical macroscopic universe with the field taking positive values initially and moving towards negative values. In this way, the considered initial conditions correspond to starting from the contracting phase of the macroscopic universe as described above. As we show in our analysis, even for not so favorable initial conditions, dynamical evolution results in ekpyrosis at a later stage in evolution causing isotropization.
@@
Recall that in the ekpyrotic phase, the energy density grows faster than anisotropies as the universe contracts towards the singularity potentially leading to its isotropization and potentially avoiding the BKL instability [48, 49]. Recently, ekpyrotic mechanisms have been used in models with ultra-slow contraction to explain homogeneity, flatness and isotropy in the contracting branch [50, 51]. Models which are ekpyrotic-like, with a different negative potential, can also potentially allow an isotropization [52, 53]. If the ekpyrotic field is combined with a bounce, such as in LQC, one can expect the bounce to isotropize, at least when one starts with anisotropies which are small. While the intuitive reason for the ekpyrotic field to diminish the growth of anisotropies in the contracting branch is easy to understand, whether or not sufficient and effective ekpyrosis occurs depends on the details of the background dynamics and is potentially sensitive to the initial conditions, including the magnitude of initial anisotropic shear compared to matter energy density. Therefore an important question with any ekpyrotic mechanism is its duration and strength and whether it can result in isotropization for arbitrary initial conditions, especially if the anisotropies are large. An example of this occurs in bouncing models in LQC where ekpyrotic potentials have been considered in both isotropic and anisotropic models showing that a non-singular evolution is possible. Answering the above question on the duration and strength of ekpyrosis is important to construct viable models of the early universe where curvature perturbations start in the contracting branch before the bounce, such as in the matter-bounce alternative to inflation [55, 56, 57, 54]. Since we consider initial conditions corresponding to the macroscopic classical contracting universe, the numerical simulations performed in this manuscript are only for the initial conditions when the field is in the positive regime and moving towards smaller $\phi$ values. This does not guarantee that ekpyrosis occurs for considered initial conditions and whether or not the universe undergoes ekpyrosis is determined by the dynamical evolution. An advantage of considering these initial conditions is two fold. First, these initial conditions are imposed in a classical macroscopic universe starting far from the quantum gravitational regime which allows to study and compare evolution of different quantum anisotropic universes with similar initial conditions. And second, by considering such initial conditions we are exploring isotropization in a more difficult setting where one does not assume favorable initial conditions for ekpyrosis. [^1] In contrast, one can always assume initial conditions which favor ekpyrosis, such as starting in the regime when field takes negative values. However, our goal in this paper is to consider initial conditions which do not start with ekpyrosis and let dynamics determine whether isotropization occurs.
@@
The goal of this manuscript is to understand whether starting from a macroscopic classical universe with an ekpyrotic potential and the scalar field rolling down from the positive values, there can be an effective isotropization of the non-singular bounce. We consider Bianchi-I and Bianchi-IX models in the effective spacetime description of LQC as our setting and contrast the dynamics when the matter-energy is sourced by a massless scalar field only versus the ekpyrotic field both in the ekpyrotic potential [47] as well as an ekpyrotic-like potential [52]. Specifically, we study whether the ekpyrotic field can dominate the bounce regime and lead to a reduction in the strength of the anisotropic shear relative to the energy density of the universe (which is assumed to be isotropic). It is important to note that we neither make any assumptions about the existence of ekpyrotic phase nor we consider specific initial conditions which favor existence of such a phase. Rather, the initial conditions we consider correspond to a contracting macroscopic universe from classical scales towards the big crunch, specifically when the field starts from positive values and rolls towards negative values in the ekpyrotic potential. The goal of this work is not to construct a viable ekpyrotic model but to understand the way ekpyrotic potential might help in isotropization. Earlier studies with ekpyrotic potential in LQC confirm that non-singular cyclic evolution is possible [58, 59], but it is only in the presence of anisotropies [60] that the moduli field can turn around which is necessary for a viable ekpyrotic scenario as proposed originally [46]. Similarly, attempts have been made to explore the effect of anisotropic evolution on the cosmological perturbations in LQC, however these studies also restrict anisotropies assuming a point-like bounce and a fixed ultra-stiff equation of state [61, 62]. But it should be noted that in general, the anisotropic approach to singularity (and hence the bounce) would be cigar-like with two scale factors contracting and one expanding before the bounce. To obtain a point-like bounce (all scale factors contracting pre-bounce) under generic anisotropic conditions it is necessary to have a viable ekpyrotic scenario. Thus, it is important to understand whether ekpyrosis can isotropize an anisotropic bounce. In this manuscript, we go beyond the assumption of small anisotropic perturbations and explore whether an isotropic universe can be obtained after bounce(s) starting from a contracting macroscopic anisotropic universe in the classical regime in the presence of an ekpyrotic potential without making any restrictions on the equation of state.
@@
The effective equations of motion for the loop quantized Bianchi-I and Bianchi-IX spacetimes are complicated and efforts are still underway to develop an intuitive understanding of the interplay between the energy density and the anisotropic shear in these models. Recently, a parabolic relation between the anisotropic shear and the energy density has been found to hold at the bounce for certain matter fields in the Bianchi-I model using extensive numerical simulations [63]. This is very surprising in light of the complicated nature of the effective equations and the fact that no such simple relation is found to hold before or after the bounce. However, the current state of the art is far from the level of analytical control needed to go beyond making a few qualitative remarks, and the progress is more often based on numerical simulations. Consequently, our analysis in this manuscript is numerical, yet wide enough to gain general insights on possible outcomes from a matter-ekpyrotic bounce. We consider the effective Bianchi-I and Bianchi-IX spacetimes for ekpyrotic and ekpyrotic-like potentials. In particular, we consider two different values of the strengths of ekpyrotic and ekpyrotic-like potentials and perform $500$ numerical simulations each with varying initial conditions chosen randomly in regime when the universe is initially classical, macroscopic and contracting towards a bounce. Since we are working in the Hamiltonian framework, the phase space variables are three triad components $p_{i}$, three connection components $c_{i}$, and the matter phase space variables: the scalar field $\phi$ and its conjugate $p_{\phi}$. Since we start the numerical simulations in the classical regime we take the initial values of the scalar field and its momentum such that the energy density $\rho$ is in the range $\rho\in[0,10^{-4}]$ in Planck units. Further, the value of $\phi$ is taken to be in the range $\phi\in[0,0.4]$. Given the initial values of $\phi$ and $\rho$, the initial value of $p_{\phi}$ (or $\dot{\phi}$) can be determined. Note that the imposition of Hamiltonian constraint reduces one degree of freedom. As a result, we provide initial conditions of three triads and two connection components only in the gravitational sector. The remaining connection component is determined from the Hamiltonian constraint. To measure the extent of the isotropization achieved due to ekpyrotic potentials, we use the massless scalar field as the reference case against which the extent of isotropization is measured. Specifically, each simulation with the ekpyrotic or ekpyrotic-like potentials is accompanied by a corresponding simulation with only the massless scalar field with same initial conditions for three triads, two connection components, energy density and the value of the massless scalar field in the range discussed above. The initial conditions for massless case are obtained by putting potential term to zero while keeping the initial conditions for three triads, two connection components, $\phi$ and $\rho$ to be same. Since the Hamiltonians of massless and non-zero potential are slightly different in the range of energy densities we perform our simulations, the initial value of connection component determined using Hamiltonian constraint for massless and non-zero potential cases is slightly different at the classical scales. Comparing the relative strength of the anisotropies at the bounce (relative to the matter energy density) we state that isotropization has been achieved if the relative strength of the anisotropies has decreased in comparison with the simulation with massless scalar field with matching initial conditions.
@@
Due to quantum geometry effects there are theoretical maximum values of energy density and anisotropic shear, given respectively by $\rho_{\mathrm{max}}\approx 0.41\rho_{\mathrm{Pl}}$ and $\sigma^{2}_{\mathrm{max}}\approx 11.57l_{\mathrm{Pl}}^{-2}$. As a result of these bounds, we find that even though the initial values of $\phi$ and $\dot{\phi}$ are not large enough to yield a long range of ekpyrosis, isotropization is achieved in an overwhelming majority of the simulations, in both the Bianchi-I and Bianchi-IX models when the ekpyrotic field is present. Even though the equation of state of the ekpyrotic fields becomes ultra-stiff only for a small duration near the bounce, isotropization can occur. We find that in ekpyrotic potentials, with the value of the field taken as positive initially, a macroscopic contracting classical universe generically undergoes multiple non-singular bounces in rapid succession before the emergence of a macroscopic expanding universe, causing multiple short spurts of ekpyrosis. These short spurts of ekpyrosis coupled with multiple bounces are found to be sufficient to produce isotropization. Another important insight we obtain is that the effect of the ekpyrotic potential on the anisotropic shear is actually indirect. This information is distilled by analyzing the effect of the ekpyrotic field on the energy density and the anisotropic shear, separately, in each simulation. We find that the ekpyrotic potential leads to a direct rise in the energy density at the bounce, indicated by a rise in the energy density at the bounce compared to the massless scalar field case in an overwhelming majority of the simulations. Since the spacetime curvature at the bounce gets contribution both from energy density and anisotropic shear, the relative strength of anisotropies in general decreases in the bounce regime. It turns out that the isotropization is achieved in a slightly higher fraction of cases in the Bianchi-I model as compared to the Bianchi-IX model. Further, we find that increasing the strength of the ekpyrotic potentials does result in a higher fraction of cases where isotropization is achieved. These findings are summarized in a table at the end of the manuscript for 4000 numerical simulations which provides a more granular look at these results.